# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ Abstract.
Reports regarding the misuse of Generative AI (GenAI) to create harmful deepfakes are emerging daily. Recently, defensive watermarking, which enables GenAI providers to hide fingerprints in their images to later use for deepfake detection, has been on the rise. Yet, its potential has not been fully explored. We present UnMarker— the first practical universal attack on defensive watermarking. Unlike existing attacks, UnMarker requires no detector feedback, no unrealistic knowledge of the scheme or similar models, and no advanced denoising pipelines that may not be available. Instead, being the product of an in-depth analysis of the watermarking paradigm revealing that robust schemes must construct their watermarks in the spectral amplitudes, UnMarker employs two novel adversarial optimizations to disrupt the spectra of watermarked images, erasing the watermarks. Evaluations against the SOTA prove its effectiveness, not only defeating traditional schemes while retaining superior quality compared to existing attacks but also breaking semantic watermarks that alter the image’s structure, reducing the best detection rate to $43\%$ and rendering them useless. To our knowledge, UnMarker is the first practical attack on semantic watermarks, which have been deemed the future of robust watermarking. UnMarker casts doubts on the very penitential of this countermeasure and exposes its paradoxical nature as designing schemes for robustness inevitably compromises other robustness aspects.
Deepfake Watermarking, Adversarial ML, Black-Box, Query-Free
ccs: Security and privacy Authenticationccs: Security and privacy Authorizationccs: Security and privacy Social network security and privacy
1. Introduction
@@
GenAI has amassed popularity due to its growing sophistication (Connor, 2023) and accessibility (Goo, [n. d.]; Cli, [n. d.]; DAL, [n. d.]). This technology has solved problems in various fields (Engelhardt et al., 2019; Prezja et al., 2022; Silbey and Hartzog, 2018; Sorin et al., 2020; Crystal et al., 2020; Lalla et al., 2022). Yet, it comes with the risk of malicious entities misusing it to generate harmful images (deepfakes) (De Ruiter, 2021; Hine and Floridi, 2022), proven by the many reports of deepfakes used in political smear campaigns (Agarwal et al., 2019) and non-consensual pornography (Harwell, 2018).
@@
Defenses first emerged as passive detectors (Chen et al., 2023; Jeong et al., 2022; Li et al., 2020; Frank et al., 2020; Zhao et al., 2021) that tell apart fake from real images, but these were proven ineffective, failing to generalize to unseen synthesis algorithms (Corvi et al., 2023). As training advanced generators is costly and prohibitive (Karras et al., 2021) in the absence of the extensive resources and high-quality data that only a few GenAI providers possess, adversaries typically resort to these providers’ APIs to generate harmful content. This recently led to the adoption of defensive watermarking as a countermeasure, which enables providers to embed hidden signatures in their images that can later be identified using designated detectors. Defensive watermarking has gained traction in the research community (Wen et al., 2023; Zhang et al., 2020; Zhu et al., 2018; Fernandez et al., 2023; Yu et al., 2022, 2021; Lukas and Kerschbaum, 2023; Tancik et al., 2020; Zhang et al., 2019; Zhao et al., 2023a) and the industry, as providers have already implemented and are testing such solutions (e.g., Google’s SynthID (Gowal, 2023)) or pledged to do so soon (Belanger, 2023) in commitments to the White House (Heikkilä;, 2023).
@@
Defensive watermarking protects against limited adversaries, following the assumption that attackers lack the resources to train complex generators. Yet, existing schemes fail to demonstrate strong guarantees for this resistance beyond empirical robustness to naive manipulations. We devise UnMarker— the first universal (effective against all schemes), black-box (without access to the scheme’s parameters or similar systems), data-free (no additional data needed) and query-free (does not rely on detector feedback) attack.
@@
Our motivating insight is that, for any scheme, there must be a universal carrier that embeds the watermark in all images. In layman’s terms, this carrier is the set of all image attributes the detector may examine when verifying the watermark. This set must exist and is also finite (shared among all images) since detectors are computationally limited. By analyzing the desired resistance to basic manipulations and the requirement that watermarks must naturally integrate with various contents, we deduce that this carrier can only be in spectral amplitudes. That is, a “conventionally” robust scheme must construct its watermarks by manipulating the frequencies at which the different pixels vary, forcing the collective distribution of pixel value shifts at different rates to follow a unique pattern. We can roughly even isolate the frequency range in which watermarks are encoded: the more robust semantic schemes (Wen et al., 2023; Tancik et al., 2020), which alter the structure, manipulate low frequencies while content-preserving non-semantic watermarks (Yu et al., 2022; Zhu et al., 2018; Lukas and Kerschbaum, 2023; Fernandez et al., 2023) populate higher bands.
@@
UnMarker thus adversarially disrupts these spectral amplitudes. To remove non-semantic watermarks, UnMarker directly modifies the image to optimize a designated loss, significantly affecting the rates at which individual pixels vary w.r.t. their neighbors, drastically affecting the distribution of high-frequency amplitudes. For semantic watermarks, we devise a novel class of optimizable filters and a procedure to learn their weights to systemically alter the consistency at different regions, causing discrepancies at lower frequencies. Unlike known filters that rely on geometric heuristics, limiting them to insignificant distortions, UnMarker is guided by a perceptual loss to maximize the disturbance in visually non-critical areas without compromising quality. As a result, the spectral carrier set is shifted away from the watermark’s distribution. By targeting the carrier, UnMarker abandons the dependence on a specific model, leaving it successful without knowledge of the scheme or detector feedback. Previous adversarial attacks (Saberi et al., 2024; Jiang et al., 2023; An et al., 2024; Lukas et al., 2024; Hu et al., 2024; Wang et al., 2022) all face this limitation, causing them to fail under realistic conditions.
@@
We experiment with seven SOTA schemes from top venues. By combining the above techniques with mild cropping ($10\%$), UnMarker breaks them all, despite its black-box and query-free nature. We compare UnMarker with the recent VAEAttack (Zhao et al., 2023c) and DiffusionAttack (Zhao et al., 2023c; Saberi et al., 2024; An et al., 2024), demonstrating far superior results, defeating semantic watermarks against which they fail and disputing misconceptions regarding these schemes’ robustness to limited adversaries (Zhao et al., 2023c; An et al., 2024; Saberi et al., 2024). Our results urge vendors to rethink their response to deepfakes. We make the following contributions:
@@
  • We thoroughly analyze defensive watermarking and reveal trade-offs that make it an invalid deepfake countermeasure.
@@
  • We present UnMarker, the first universally-effective data-free, black-box, and query-free attack against defensive watermarking [^1] Our code will be made publicly available upon publication. UnMarker’s success is due to its novel spectral optimizations.
@@
  • We demonstrate UnMarker’s efficacy against seven SOTA schemes under restrictive settings, rendering them of no practical use.
@@
  • We compare UnMarker to the most powerful black-box and query-free attacks, showing superior performance and image quality.
@@
  • Using a novel class of optimizable filters to incur structural disturbances, UnMarker is the first practical attack to defeat semantic watermarks, whose previous unparalleled robustness has marked them as the ultimate candidates for future trustworthy deepfake detection, casting doubts on the feasibility of this task.
2. Background & Related Work
Here, we provide related work on deepfake detection and defensive watermarking. We survey image synthesis techniques in A.1.
2.1. Deepfake Detection
@@
Passive Detectors. They distinguish between real and fake images (Chen et al., 2023; Jeong et al., 2022; Li et al., 2020; Frank et al., 2020; Zhao et al., 2021) based on semantic incoherences (Hu et al., 2021; Matern et al., 2019) and low-level artifacts (Chen et al., 2023; Jeong et al., 2022; Frank et al., 2020; Qian et al., 2020; Marra et al., 2019; Wu et al., 2020) in deepfakes due to computational limitations. As proven in the image (Dong et al., 2022; Cao and Gong, 2021) and audio domains (Kassis and Hengartner, 2023), advanced generators can produce high-quality outputs without such telltales, and post-processing manipulations can fool these detectors that also fail to generalize to new systems (Corvi et al., 2023).
@@
Defensive Watermarking. It assumes limited attackers compared to GenAI providers, lacking the resources or data to train high-quality generators. A watermarking scheme consists of watermark ($w$), encoder ($\mathcal{E}$) and detector ($\mathcal{D}$). In a general-purpose scheme (Zhang et al., 2020; Zhu et al., 2018; Tancik et al., 2020), $\mathcal{E}$ accepts an arbitrary image $x$ and outputs a similar watermarked $x_{w}=\mathcal{E}(x,w)$. Given image $x^{\prime}$, $\mathcal{D}$ extracts a sequence $w^{\prime}=\mathcal{D}(x^{\prime})$ that will be almost identical to $w$ $iff$ $x^{\prime}$ was previously watermarked. Generator-specific approaches “watermark” generative models: The encoder $\mathcal{E}$ can be the generator $\mathcal{G}$ itself that is trained to only produce watermarked images (Yu et al., 2022, 2021) or an algorithm that modifies $\mathcal{G}$’s weights (Lukas and Kerschbaum, 2023; Fernandez et al., 2023) or latent codes (Wen et al., 2023) s.t. every image it produces embeds the watermark. Heuristics-based watermarks (Bi et al., 2007; Pereira and Pun, 2000; ShieldMnt, 2021; Wen et al., 2023) directly manipulate spectral representations (Bi et al., 2007; Pereira and Pun, 2000; ShieldMnt, 2021), forming a signature in select frequencies. These are removable via post-processing such as JPEG (Jiang et al., 2023; Fernandez et al., 2023; Zhu et al., 2018). We focus on learning-based methods (Zhang et al., 2020; Zhu et al., 2018; Tancik et al., 2020; Yu et al., 2022, 2021; Lukas and Kerschbaum, 2023; Fernandez et al., 2023; Wen et al., 2023; Zhang et al., 2019; Zhao et al., 2023b), where $\mathcal{E}$ and $\mathcal{D}$ are DNNs as they are robust to post-processing.
Semantic vs. Non-Semantic: Non-semantic schemes take an image and produce a visually identical watermarked version. Aside from StegaStamp (Tancik et al., 2020), all general-purpose schemes (Zhang et al., 2020; Zhu et al., 2018; Zhang et al., 2019) are non-semantic. Generator-specific schemes (Yu et al., 2022, 2021; Lukas and Kerschbaum, 2023; Fernandez et al., 2023), with the exception of TRW (Wen et al., 2023), are also non-semantic, as their training restricts them to output images identical to non-watermarked references.
Semantic watermarks can influence the content. This term was popularized by Wen et al. with their introduction of TRW (Wen et al., 2023) that alters the latent codes used by conditional text-to-image Latent Diffusion Models (LDMs) (Rombach et al., 2022). When providing a “random” code with an auxiliary phrase that controls the output, the LDM uses this code as a seed to generate an image matching the text (e.g, the phrase “a dog on a bike” leads the LDM to output an image of a dog riding a bike but the dog and bike will differ based on the seed). TRW injects a universal pattern in the Fourier transforms of all latent codes before generation. As a result, watermarked images will be different from non-watermarked ones but still satisfy the text. This leaves watermarks invisible despite being highly robust.
@@
StegaStamp (Tancik et al., 2020) precedes TRW and operates similar to its general-purpose counterparts. Yet, it is less stealthy, causing structural changes (see Fig 2). Still, its strategically distributed distortion ensures minimal visibility. In defensive watermarking, users cannot observe non-watermarked images, further hiding these nuances and making StegaStamp highly applicable. While semantic watermarks primarily refer to TRW, previous work recognizes the similarities in both schemes, distinguishing them as more robust (Zhao et al., 2023c; Saberi et al., 2024; An et al., 2024). Thus, we refer to TRW and StegaStamp as semantic.
2.2. Attacks Against Defensive Watermarking
@@
Regeneration Attacks. Initially proposed by Zhao et al. (Zhao et al., 2023c), they rely on Diffusion Models (DMs) (Sohl-Dickstein et al., 2015) (see §A.1) or Variational Autoencoders (Kingma and Welling, 2013) to introduce noise and then reconstruct the images, restoring their quality after the watermark has been eliminated. Saberi et al. (Saberi et al., 2024) explore similar attacks, and WAVES (An et al., 2024) repeatedly propagates images through the regeneration pipeline to maximize the disruption at the risk of decreasing the sample’s quality. All three works verify the vulnerability of non-semantic watermarks against which regeneration attacks offer theoretical guarantees. Yet, semantic schemes remain robust. While WAVES shows that TRW experiences some degradation, it only occurs at a considerable distortion, at which point the attack loses its validity (see §3).
Regeneration methods remain the only practical (black-box and query-free) attacks, but they fail against semantic watermarks. Their success, even against non-semantic schemes, also depends on the denoising pipelines’ fidelity (due to computational power and training data), often severely impacting quality (see §6.2.3). UnMarker is the first to defeat all watermarks while also abandoning the dependence on regeneration models, always ensuring quality.
@@
Adversarial Attacks. Practical attacks must assume no access to the schemes’ internals (see §3). This makes traditional adversarial attacks that target the decision boundaries of a specific model ineffective. Thus, existing works make relaxing assumptions, failing to retain success under realistic conditions.
WEVADE (Jiang et al., 2023) optimizes watermarked images by querying the detector and updating them accordingly until the watermark is no longer detected. However, detector feedback is not necessarily available (see §3), making it ineffective. Lukas et al. (Lukas et al., 2024) propose powerful query-free adversarial attacks, but they assume unrealistic access to a generator with similar architecture and data (see §3).
Saberi et al. (Saberi et al., 2024) propose a surrogate-based attack. The scheme is defeated by mounting an adversarial attack against the surrogate, which transfers to the target. Training the substitute requires non-watermarked samples from the provider’s distribution, violating the principles of defensive watermarking (An et al., 2024; Hu et al., 2024) (see §3) as providers only release watermarked images and possess exclusive datasets. Still, the attack fails against the semantic StegaStamp without considerably affecting quality. RD-IWAN (Wang et al., 2022) shares similar pitfalls, requiring access to non-watermarked provider images.
WAVES (An et al., 2024) evaluates surrogate attacks specifically on semantic schemes: StegaStamp remains robust, while TRW is only affected under unrealistic access to its autoencoder or images generated by it with different watermarks, which is impractical as a provider only publishes images with its single identifying watermark. They also show how Saberi et al.’s (Saberi et al., 2024) surrogate attack fails even against non-semantic schemes absent any non-watermarked provider images.
Hu et al.’s (Hu et al., 2024) ensemble surrogate attack requires no non-water-marked provider images. Yet, evaluations are restricted to surrogates and targets of the same scheme (training methods and objectives), leading to sufficient commonalities for transferability and violating the black-box setting. It also requires extensive resources (training 40-100 surrogates), data, and large perturbations, making it impractical. The paper does not consider semantic watermarks.
UnMarker waives all impractical assumptions, requires no data or prohibitive computational power (see §7), and preserves quality.
3. Threat Model

### v2

# UnMarker: A Universal Attack on Defensive Image Watermarking

@@ Abstract
Reports regarding the misuse of Generative AI (GenAI) to create deepfakes are frequent. Defensive watermarking enables GenAI providers to hide fingerprints in their images and use them later for deepfake detection. Yet, its potential has not been fully explored. We present UnMarker— the first practical universal attack on defensive watermarking. Unlike existing attacks, UnMarker requires no detector feedback, no unrealistic knowledge of the watermarking scheme or similar models, and no advanced denoising pipelines that may not be available. Instead, being the product of an in-depth analysis of the watermarking paradigm revealing that robust schemes must construct their watermarks in the spectral amplitudes, UnMarker employs two novel adversarial optimizations to disrupt the spectra of watermarked images, erasing the watermarks. Evaluations against SOTA schemes prove UnMarker’s effectiveness. It not only defeats traditional schemes while retaining superior quality compared to existing attacks but also breaks semantic watermarks that alter an image’s structure, reducing the best detection rate to $43\%$ and rendering them useless. To our knowledge, UnMarker is the first practical attack on semantic watermarks, which have been deemed the future of defensive watermarking. Our findings show that defensive watermarking is not a viable defense against deepfakes, and we urge the community to explore alternatives.
1 Introduction
@@
GenAI has amassed popularity due to its growing sophistication and accessibility [1, 2, 3]. This technology has solved problems in various fields [4, 5, 6, 7]. Yet, it comes with the risk of malicious entities using it to generate harmful images (deepfakes) [8, 9], proven by many reports of those being used in political smear campaigns [10] and non-consensual pornography [11].
@@
Defenses first emerged as passive detectors [12, 13, 14, 15] that tell apart fake from real images, but these were proven ineffective, failing to generalize to unseen synthesis algorithms [16]. As training advanced generators is costly and prohibitive [17] in the absence of the extensive resources and high-quality data that only a few GenAI providers possess, adversaries typically resort to these providers’ APIs to generate harmful content. This recently led to the adoption of defensive watermarking as a countermeasure, which enables providers to embed hidden signatures in their images that can later be identified using designated detectors. Defensive watermarking has gained traction in the research community [18, 19, 20, 21, 22, 23, 24, 25] and the industry, as providers have already implemented and are testing such solutions (e.g., Google’s SynthID [26]) or pledged to do so soon [27] in commitments to the White House [28].
@@
Defensive watermarking protects against limited adversaries, following the assumption that attackers lack the resources to train complex generators. Yet, existing schemes fail to demonstrate strong guarantees for this resistance beyond empirical robustness to naive manipulations. We devise UnMarker— the first universal (effective against all schemes), black-box (without access to the scheme’s parameters or similar systems), data-free (no additional data needed) and query-free (without detector feedback) attack.
@@
Our motivating insight is that, for any scheme, there must be a universal carrier that embeds the watermark in all images. In layman’s terms, this carrier is the set of all attributes the detector may examine when verifying the watermark. This set must exist and is finite (shared among all images) as detectors are computationally limited. By analyzing the desired resistance to basic manipulations and the requirement that watermarks must naturally integrate with various contents, we deduce that this carrier can only be in spectral amplitudes. That is, a “conventionally” robust scheme must construct its watermarks by manipulating the frequencies at which the different pixels vary, forcing the collective distribution of these value shifts at different rates to follow a unique pattern. We can roughly even isolate the range in which watermarks are encoded: the more robust semantic schemes [18, 24], which alter the structure, target low frequencies while content-preserving non-semantic watermarks [21, 19, 23, 20] populate higher bands.
@@
UnMarker disrupts these spectral amplitudes. To remove non-semantic watermarks, it directly modifies the image to optimize a designated loss, affecting the rates at which individual pixels vary w.r.t. their neighbors and impacting the distribution of high-frequency amplitudes. For semantic watermarks, we devise novel optimizable filters and a procedure to learn their weights to systemically alter the consistency at different regions, causing discrepancies at lower frequencies. Unlike known filters that rely on geometric heuristics, limiting them to minor distortions, UnMarker is guided by a perceptual loss to maximize the disturbance in visually non-critical areas without compromising quality. As a result, the carrier set is shifted away from the watermark’s distribution. By targeting the carrier, UnMarker abandons the dependence on a specific model, leaving it successful without knowledge of the scheme or detector feedback. Previous adversarial attacks [29, 30, 31, 32, 33, 34] all face this limitation, failing under realistic conditions.
@@
We experiment with seven SOTA schemes from top venues. By combining the above techniques with mild cropping, UnMarker breaks them all, despite its black-box and query-free nature. We compare UnMarker with the recent VAEAttack [35] and DiffusionAttack [35, 29, 31], demonstrating far superior results, defeating semantic watermarks against which they fail and disputing misconceptions regarding these schemes’ robustness to limited adversaries [35, 31, 29]. Our results urge vendors to rethink their response to deepfakes. We make the following contributions:
@@
  • We thoroughly analyze defensive watermarking and reveal trade-offs that make it an invalid countermeasure.
@@
  • We employ novel spectral optimizations to devise UnMarker— the first universal data-free, black-box, and query-free attack against defensive watermarking [^1] https://github.com/andrekassis/ai-watermark.
@@
  • We demonstrate UnMarker’s efficacy against seven SOTA schemes under restrictive settings, rendering them useless.
@@
  • We compare UnMarker to the strongest query-free black-box attacks, attaining superior performance and quality.
@@
  • Using a novel class of optimizable filters to incur structural disturbances, UnMarker is the first practical attack to defeat semantic watermarks, whose robustness has marked them as the ultimate candidates for trustworthy deepfake detection, showing this assumption is highly questionable.
2 Background & Related Work
2.1 Deepfake Detection
@@
Passive Detectors. These systems [12, 13, 14, 15] identify deepfakes based on semantic incoherences [36, 37] and low-level artifacts [12, 14, 38, 39]. Advanced generators can produce high-quality outputs without such telltales, and manipulations can fool these detectors [40, 41, 42]. They also fail to generalize to unseen models [16].
@@
Defensive Watermarking. It assumes attackers lack the resources to train high-quality generators. A watermarking scheme consists of watermark ($w$), encoder ($\mathcal{E}$) and detector ($\mathcal{D}$). In a general-purpose scheme [19, 24], $\mathcal{E}$ accepts any image $x$ and outputs a similar watermarked $x_{w}\!\!=\!\!\mathcal{E}(x,w)$. Given image $x^{\prime}$, $\mathcal{D}$ extracts a sequence $w^{\prime}\!\!=\!\!\mathcal{D}(x^{\prime})$ that will be almost identical to $w$ $iff$ $x^{\prime}$ was previously watermarked. Generator-specific approaches “watermark” generative models: $\mathcal{E}$ can be the generator itself trained to only produce watermarked images [21, 22] or an algorithm that modifies its weights [23, 20] or latent codes [18] s.t. images it produces embed the watermark. Heuristics-based watermarks [43, 44, 45] directly manipulate spectral representations, forming a signature in select frequencies. These are removable via post-processing [30, 20]. We focus on learning-based methods [19, 24, 21, 22, 23, 20, 18, 25], where $\mathcal{E}$ and $\mathcal{D}$ are DNNs as they resist post-processing.
@@
Semantic vs. Non-Semantic: Non-semantic schemes take an image and produce a visually identical watermarked version. Aside from StegaStamp [24], general-purpose schemes are non-semantic. Generator-specific schemes, except TRW [18], are also non-semantic, as their training restricts them to output images identical to non-watermarked references. Semantic watermarks can influence the content. This term was popularized by Wen et al. with their introduction of TRW [18] that alters the latent codes used by text-to-image Latent Diffusion Models (LDMs) [46]. When providing a random code with a textual phrase, the LDM uses this code as a seed to produce an image matching the text (e.g, “a dog on a bike” results in an image of a dog riding a bike but these objects will change based on the seed). TRW injects a pattern in the Fourier transforms of all latent codes before generation. As a result, watermarked images will differ from non-watermarked ones but still satisfy the text. This leaves watermarks invisible but highly robust. StegaStamp [24] precedes TRW and operates similar to its general-purpose counterparts. Yet, it is less stealthy, causing structural changes (see Fig 2). Still, its strategically distributed distortion ensures minimal visibility. In defensive watermarking, users cannot observe non-watermarked images, further hiding these nuances and making StegaStamp highly applicable. While semantic watermarks primarily refer to TRW, previous work recognizes the similarities in both schemes, distinguishing them as more robust [35, 29, 31]. Thus, we refer to TRW and StegaStamp as semantic.
2.2 Attacks Against Defensive Watermarking
@@
Regeneration Attacks. Proposed by Zhao et al. [35], they rely on Diffusion Models (DMs) [47] or Variational Autoencoders (VAEs) [48] (henceforth, the DiffusionAttack and VAEAttack) to introduce noise and then reconstruct images after the watermark has been eliminated. Saberi et al. [29] and WAVES [31] revisit these attacks. All these works verify non-semantic watermarks’ vulnerability, against which regeneration attacks offer theoretical guarantees. Yet, semantic schemes remain robust. While TRW experiences some degradation [31], it occurs at a considerable distortion where the attack is invalid (see §3). Regeneration is the only practical (black-box and query-free) attack but fails against semantic watermarks. It also relies on the denoising pipelines’ fidelity, often impacting quality (see §6.2.3). UnMarker defeats all watermarks while also abandoning the dependence on regeneration models, ensuring quality.
@@
Adversarial Attacks. Realistically, attackers cannot access the scheme (see §3), making traditional adversarial attacks that target the decision boundaries of specific models ineffective. Thus, existing works make relaxing assumptions, failing to retain success under practical conditions.
WEVADE [30] attacks watermarked images by querying the detector and updating them accordingly until the watermark is not detected. Yet, detector feedback may not be available (see §3), making it ineffective. Lukas et al. [32] propose query-free attacks but assume unrealistic access to a generator with similar architecture and data (see §3). Saberi et al. [29] mount an attack on a surrogate detector, which transfers to the target. Training the substitute requires non-watermarked samples from the provider’s distribution, violating the defensive watermarking principles [31, 33] (see §3) as providers only release watermarked images and possess exclusive datasets. Still, it fails against the semantic StegaStamp without affecting quality. RD-IWAN [34] also requires access to non-watermarked provider images. An et al. [31] evaluate surrogate attacks specifically on semantic schemes, finding StegaStamp robust, while TRW is affected under unrealistic access to its autoencoder or images generated by it with different watermarks, which is impractical as any provider’s published image bear its single identifying watermark. They also prove Saberi et al.’s [29] surrogate attack unsuccessful absent non-watermarked provider images. Hu et al.’s [33] ensemble surrogate attack requires no non-watermarked provider images. Yet, the evaluations exclude semantic watermarks and only consider surrogates and targets of the same scheme, leading to sufficient commonalities for transferability and violating the black-box setting. It also requires extensive resources ($40\!\!-\!\!100$ surrogates), data, and large perturbations, making it impractical. UnMarker waives impractical assumptions, requires no data or prohibitive resources, and preserves quality.
3 Threat Model
Our framework considers two parties: Providers who develop and monetize these platforms, and Attackers abusing their access to generate malicious content.